A truck travels at a speed of 80.0 km/h and has a mass of 2 000 kg. What is its kinetic energy?
1
Expert's answer
2020-06-10T18:23:00-0400

By the definition of the kinetic energy, we get:


"KE = \dfrac{1}{2}mv^2,""KE = \dfrac{2000 \ kg}{2} \cdot (80 \ \dfrac{km}{h} \cdot \dfrac{1000 \ m}{1 \ km} \cdot \dfrac{1 \ h}{3600 \ s})^2 = 493827 \ J."

Answer:

"KE = 493827 \ J."